Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
4x – 8y = 48
11x + 3y = -105.5

Found 2 solutions by richwmiller, mananth:
Answer by richwmiller(17219)

x~~-7., y~~-9.5

Answer by mananth(16949)

4 x -8 y = 48 .............1

11 x + 3 y = -105.50 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 8
12 x -24 y = 144
88 x + 24 y = -844
Add the two equations
100 x = -700
/ 100
x = -7
plug value of x in (1)
4 x -8 y = 48
-28 -8 y = 48
-8 y = 48 + 28
-8 y = 76
y = -9 .5
x=-7, y = -9.5